# GI-NAS: Boosting Gradient Inversion Attacks through Adaptive Neural Architecture Search

### v2

@@ Abstract
Gradient Inversion Attacks invert the transmitted gradients in Federated Learning (FL) systems to reconstruct the sensitive data of local clients and have raised considerable privacy concerns. A majority of gradient inversion methods rely heavily on explicit prior knowledge (e.g., a well pre-trained generative model), which is often unavailable in realistic scenarios. To alleviate this issue, researchers have proposed to leverage the implicit prior knowledge of an over-parameterized network. However, they only utilize a fixed neural architecture for all the attack settings. This would hinder the adaptive use of implicit architectural priors and consequently limit the generalizability. In this paper, we further exploit such implicit prior knowledge by proposing Gradient Inversion via Neural Architecture Search (GI-NAS), which adaptively searches the network and captures the implicit priors behind neural architectures. Extensive experiments verify that our proposed GI-NAS can achieve superior attack performance compared to state-of-the-art gradient inversion methods, even under more practical settings with high-resolution images, large-sized batches, and advanced defense strategies.
@@ I Introduction
Federated Learning (FL) [1, 2, 3, 4] serves as an efficient collaborative learning framework where multiple participants cooperatively train a global model and only the computed gradients are exchanged. By adopting this distributed paradigm, FL systems fully leverage the huge amounts of data partitioned across various clients for enhanced model efficacy and tackle the separateness of data silos [5]. Moreover, since merely the gradients instead of the intimate data are uploaded to the server, the user privacy seems to be safely guaranteed as the private data is only available at the client side.
@@
However, FL systems are actually not so reliable as what people have expected. Extensive studies have discovered that even the conveyed gradients can disclose the sensitive information of users. Early works [6, 7, 8, 9] involve inferring the existence of certain samples in the dataset (i.e., Membership Inference Attacks [10]) or further revealing some properties of the private training set (i.e., Property Inference Attacks [11]) from the uploaded gradients. But unlike the above inference attacks that only partially reveal limited information of the private data, Gradient Inversion Attacks stand out as a more threatening privacy risk as they can completely reconstruct the sensitive data by inverting the gradients.
@@
Zhu et al. [12] first formulate gradient inversion as an optimization problem and recover the intimate training images by minimizing the gradient matching loss (i.e., the distance between the dummy gradients and the real gradients) with image pixels regarded as trainable parameters. Ensuing works improve the attack performance on the basis of Zhu et al. [12] by designing label extraction techniques [13], switching the distance metric and introducing regularization [14], or considering larger batch sizes settings [15], but still restrict their optimizations in the pixel space. To fill this gap, recent studies [16, 17, 18] propose to explore various search algorithms within the Generative Adversarial Networks (GAN) [19, 20] to leverage the rich prior knowledge encoded in the pre-trained generative models. While incorporating these explicit priors indeed improves the attack performance, it is usually tough to pre-prepare such prerequisites in realistic FL scenarios where the data distribution at the client side is likely to be complex or unknown to the attackers. Therefore, Zhang et al. [21] propose to employ an over-parameterized convolutional network for gradient inversion and directly optimize the excessive network parameters on a fixed input, which does not require any explicit prior information but still outperforms GAN-based attacks. The reason behind this is that the structure of a convolutional network naturally captures image statics prior to any learning [22] and Zhang et al. [21] leverage this characteristic as implicit prior knowledge. However, Zhang et al. [21] only utilize a fixed network for all the attack settings, regardless of the specific batch to recover. An intuitive question naturally arises: Can we adaptively search the architecture of the over-parameterized network to better capture the implicit prior knowledge for each reconstructed batch?
@@
In Figure 1, we randomly select $5$ over-parameterized networks with different architectures to attack $3$ different batches. All the networks are optimized for the same number of iterations on ImageNet [23]. The results indicate that the PSNR performance varies significantly when changing the architectures. For the same batch, various architectures can hold remarkably different implicit priors on it. Besides, since the optimal models for Batch $1$, Batch $2$, and Batch $3$ are respectively Model $1$, Model $3$, and Model $4$, there exists no network that can consistently perform the best on all the given batches. Thus, the changeless adoption of architecture by Zhang et al. [21] lacks optimality under dynamic FL scenarios where different batches require different implicit architectural priors, and it is of great significance to adaptively select the most suitable architecture for each batch.
@@
  • We provide deeper analysis on various aspects, including time costs, robustness to network parameters initialization, robustness to latent codes initialization, implications behind searched results, ablation study on each search type, and generalizability to more FL global models.
@@ II-A Gradient Inversion Attacks and Defenses
Zhu et al. [12] first propose to restore the private samples via iterative optimization for pixel-level reconstruction, yet limited to low-resolution and single images. Geiping et al. [14] empirically decompose the gradient vector by its norm magnitude and updating direction, and succeed on high-resolution ImageNet [23] through the angle-based loss design. Furthermore, Yin et al. [15] extend the attacks to the batch-level inversion through the group consistency regularization and the improved batch label inference algorithm [13]. With strong handcrafted explicit priors (e.g., fidelity regularization, BN statistics), they accurately realize batch-level reconstruction with detailed semantic features allocated to all the individual images in a batch. Subsequent studies [16, 17, 18] leverage pre-trained GAN models as generative priors [24, 25, 26] to enhance the attacks. Current defenses focus on gradient perturbation to alleviate the impact of gradient inversion with degraded gradients [27, 28, 29]. Gaussian Noise [30] and Gradient Clipping [31] are common techniques in Differential Privacy (DP) [32, 33] that effectively constrain the attackers from learning through the released gradients. Gradient Sparsification [34, 35] prunes the gradients through a given threshold, and Soteria [36] edits the gradients from the aspect of learned representations.
@@ II-B Neural Architecture Search (NAS)
By automatically searching the optimal model architecture, NAS algorithms prove to be significantly effective in multiple visual tasks such as image restoration [37, 38], semantic segmentation [39, 40, 41], and image classification [42, 43, 44, 45, 46]. In image classification tasks, Zoph et al. [43, 45] regularize the search space to a convolutional cell and construct a better architecture stacked by these cells using an RNN [47] controller. For semantic segmentation tasks, Liu et al. [40] search for the optimal network on the level of hierarchical network architecture and extend NAS to dense image prediction. In terms of image restoration tasks, Suganuma et al. [37] exploit a better Convolutional Autoencoders (CAE) with standard network components through the evolutionary search, while Chu et al. [38] discover a competitive lightweight model via both micro and macro architecture search.
Existing NAS methods adopt different strategies to exploit the search space, including evolutionary methods [48, 49, 50, 51, 52, 53, 54], Bayesian optimization [55, 56, 57], Reinforcement Learning (RL) [58, 59, 60, 43, 45], and gradient-based search [61, 62, 63, 64]. Different RL-based methods vary in the way to represent the agent’s policy and the optimization process. Zoph et al. [43] utilize RNN networks to sequentially encode the neural architecture and train them with the REINFORCE policy gradient algorithm [65]. Baker et al. [58] adopt Q-learning [66, 67] to train the policy network and realize competitive model design. Notably, recent works [68, 69, 70] propose training-free NAS to mitigate the issue of huge computational expenses. Instead of training from scratch, they evaluate the searched networks by some empirically designed metrics that reflect model effectiveness. In this paper, we adopt the initial gradient matching loss as the training-free search metric and provide substantial empirical evidence that such a search metric highly correlates with the real attack performance.
@@ III Problem Formulation
GAN-based Gradient Inversion. Nevertheless, the optimization of (3) is still limited in the pixel space. Given a well pre-pretrained GAN, an instinctive idea is to swift the optimization from the pixel space to the GAN latent space:
@@
where $G_{\omega}$ and $\mathbf{z}\in\mathbb{R}^{B\times l}$ are respectively the generator and the latent vector of the pre-trained GAN. By reducing the optimization space from $\mathbb{R}^{B\times H\times W\times C}$ to $\mathbb{R}^{B\times l}$, (4) overcomes the uncertainty of directly optimizing the extensive pixels and exploits the abundant prior knowledge encoded in the pre-trained GAN. Based on this, recently emerged GAN-based attacks [16, 17] explore various search strategies within the pre-trained GAN to utilize its expression ability.
@@
where $G_{over}$ is the over-parameterized convolutional network with excessive parameters $\mathbf{\phi}$, and $\mathbf{z_{0}}$ is the randomly generated but fixed latent code. Note that the regularization term is omitted in (5). This is because the architecture of $G_{over}$ itself can serve as implicit regularization, for convolutional networks have been proven to possess implicit priors that prioritize clean images rather than noise as shown in [22]. Thus, the generated images that highly resemble the ground truth images can be obtained through $\mathbf{\hat{x}^{*}}=G_{over}(\mathbf{z_{0}};\mathbf{\phi^{*}})$. However, only a changeless over-parameterized network is employed for all the attack settings in [21]. As previously shown in Figure 1, although the network is over-parameterized, the attack performance exhibits significant differences when adopting different architectures. To fill this gap, we propose to further exploit such implicit architectural priors by searching the optimal over-parameterized network $G_{opt}$ for each batch. We will discuss how to realize this in Section IV.
@@ IV Method
Figure 2: Overview of the proposed GI-NAS attack. We leverage a two-stage strategy for private batch recovery. In the first stage, we traverse the model search space and calculate the initial gradient matching loss (i.e., our training-free metric) of each model based on the fixed input $\mathbf{z_{0}}$. We regard the model that achieves the minimal initial loss as our best model, for its performance at the start can stand from numerous candidates. In the second stage, we adopt the architecture of the previously found best model and optimize its excessive parameters to reconstruct the private data.
@@
IV-A Training-free Optimal Model Search
IV-A1 Model Search Space Design
@@
Search Space for Skip Connection Patterns. We assume that there are $t$ levels of encoders and decoders in total, and denote them as $e_{1},e_{2},...,e_{t}$ and $d_{1},d_{2},...,d_{t}$. As shown in Figure 3, We consider different skip connection patterns between encoders and decoders. To represent each of these patterns, we define a skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$ that serves as a mask to determine whether there will be new residual connections [75] between pairs of encoders and decoders. More concretely, $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there isn’t such a skip connection. As the shapes of feature maps across different network levels can vary significantly (e.g., $64\times 64$ for the output of $e_{i}$ and $256\times 256$ for the input of $d_{j}$), we introduce connection scale factors to tackle this inconsistency and decompose all the possible scale factors into a series of 2×2$\times$2 × upsampling operations or downsampling operations with shared weights. By allocating 00 or $1$ to each of the $t^{2}$ bits in A, we broaden the search space of skip connection patterns as there are $2^{t^{2}}$ possibilities altogether and we only need to sample a portion of them.
Figure 3: The design of search space for skip connection patterns. Different skip connection patterns are determined by the skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$. $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there isn’t such a skip connection.
IV-A2 Optimal Model Selection

### v3

@@ Abstract
Gradient Inversion Attacks invert the transmitted gradients in Federated Learning (FL) systems to reconstruct the sensitive data of local clients and have raised considerable privacy concerns. A majority of gradient inversion methods rely heavily on explicit prior knowledge (e.g., a well pre-trained generative model), which is often unavailable in realistic scenarios. This is because real-world client data distributions are often highly heterogeneous, domain-specific, and unavailable to attackers, making it impractical for attackers to obtain perfectly matched pre-trained models, which inevitably suffer from fundamental distribution shifts relative to target private data. To alleviate this issue, researchers have proposed to leverage the implicit prior knowledge of an over-parameterized network. However, they only utilize a fixed neural architecture for all the attack settings. This would hinder the adaptive use of implicit architectural priors and consequently limit the generalizability. In this paper, we further exploit such implicit prior knowledge by proposing Gradient Inversion via Neural Architecture Search (GI-NAS), which adaptively searches the network and captures the implicit priors behind neural architectures. Extensive experiments verify that our proposed GI-NAS can achieve superior attack performance compared to state-of-the-art gradient inversion methods, even under more practical settings with high-resolution images, large-sized batches, and advanced defense strategies. To the best of our knowledge, we are the first to successfully introduce NAS to the gradient inversion community. We believe that this work exposes critical vulnerabilities in real-world federated learning by demonstrating high-fidelity reconstruction of sensitive data without requiring domain-specific priors, forcing urgent reassessment of FL privacy safeguards.
@@ I Introduction
Federated Learning (FL) [1, 2, 3] serves as an efficient collaborative learning framework where multiple participants cooperatively train a global model and only the computed gradients are exchanged. By adopting this distributed paradigm, FL systems fully leverage the huge amounts of data partitioned across various clients for enhanced model efficacy and tackle the separateness of data silos [4]. Moreover, since merely the gradients instead of the private data are uploaded to the server, the user privacy seems to be safely guaranteed as the private data is only available at the client side.
@@
However, FL systems are actually not so secure as what people have expected [5, 6, 7, 8, 9, 10, 11, 12, 13]. Extensive studies have discovered that even the transmitted gradients can disclose the sensitive information of users. Early works [5, 6, 7, 8] involve inferring the existence of certain samples in the dataset (i.e., Membership Inference Attacks [14]) or further revealing some properties of the private training set (i.e., Property Inference Attacks [15]) from the uploaded gradients. But unlike the above inference attacks that only partially reveal limited information of the private data, Gradient Inversion Attacks [9, 10, 11, 12, 13] stand out as a more threatening privacy risk as they can completely reconstruct the sensitive data by inverting the gradients.
@@
Zhu et al. [9] first formulate gradient inversion as an optimization problem and recover the private training images by minimizing the gradient matching loss (i.e., the distance between the dummy gradients and the real gradients) with image pixels regarded as trainable parameters. Ensuing works improve the attack performance on the basis of Zhu et al. [9] by designing label extraction techniques [16], switching the distance metric and introducing regularization [10], or considering settings with larger batch sizes [17], but still restrict their optimizations in the pixel space. To fill this gap, recent studies [11, 18, 12] propose to explore various search algorithms within the Generative Adversarial Networks (GAN) [19, 20] to leverage the rich prior knowledge encoded in the pre-trained generative models. While incorporating these explicit priors indeed improves the attack performance, it is usually tough to pre-prepare such prerequisites in realistic FL scenarios where the data distribution at the client side is likely to be complex or unknown to the attackers. Therefore, Zhang et al. [13] propose to employ an over-parameterized convolutional network for gradient inversion and directly optimize the excessive network parameters on a fixed input, which does not require any explicit prior information but still outperforms GAN-based attacks. The reason behind this is that the structure of a convolutional network naturally captures image statistics prior to any learning [21] and Zhang et al. [13] leverage this characteristic as implicit prior knowledge. However, Zhang et al. [13] only utilize a fixed network for all the attack settings, regardless of the specific batch to recover. An intuitive question naturally arises: Can we adaptively search the architecture of the over-parameterized network to better capture the implicit prior knowledge for each reconstructed batch?
@@
In Figure 1, we randomly select $5$ over-parameterized networks with different architectures to attack $3$ different batches. All the networks are optimized for the same number of iterations on ImageNet [22]. The results indicate that the Peak Signal-to-Noise Ratio (PSNR) performance varies significantly when changing the architectures. For the same batch, various architectures can hold remarkably different implicit priors on it. Besides, since the optimal models for Batch $1$, Batch $2$, and Batch $3$ are respectively Model $1$, Model $3$, and Model $4$, there exists no network that can consistently perform the best on all the given batches. Thus, the rigid use of a single architecture by Zhang et al. [13] lacks optimality under dynamic FL scenarios where different batches require different implicit architectural priors, and it is of great significance to adaptively select the most suitable architecture for each batch.
@@
  • We provide deeper analysis on various aspects, such as computational efficiency, ablation studies, generalizability to more FL global models, robustness to network parameters and latent codes initialization, and implications behind the NAS searched results.
Research Purpose and Real-World Significance. This research reveals a critical security vulnerability in federated learning by developing GI-NAS, an adaptive gradient inversion attack that dynamically adjusts to diverse real-world conditions. Unlike existing methods constrained by fixed architectures or unrealistic data assumptions, our approach automatically identifies optimal network configurations through neural architecture search, enabling more effective privacy attacks across varying batch characteristics. The practical significance lies in exposing security risks in actual FL deployments where data resolutions, batch sizes, and defense mechanisms constantly vary. By demonstrating how the architectural adaptation enhances the attack effectiveness, we believe that this work will not only advance attack methodologies but also provide crucial insights for developing stronger defense systems in real-world distributed learning environments where sensitive data protection remains paramount.
@@ II-A Gradient Inversion Attacks and Defenses
Zhu et al. [9] first propose to restore the private samples via iterative optimization for pixel-level reconstruction, yet limited to low-resolution and single images. Geiping et al. [10] empirically decompose the gradient vector by its norm magnitude and updating direction, and succeed on high-resolution ImageNet [22] through the angle-based loss design. Furthermore, Yin et al. [17] extend the attacks to the batch-level inversion through the group consistency regularization and the improved batch label inference algorithm [16]. With strong handcrafted explicit priors (e.g., fidelity regularization, BN statistics), they accurately realize batch-level reconstruction with detailed semantic features allocated to all the individual images in a batch. Subsequent studies [11, 18, 12] leverage pre-trained GAN models as generative priors [23, 24, 25, 26, 27, 28, 29] to enhance the attacks. Current defenses focus on gradient perturbation to alleviate the impact of gradient inversion with degraded gradients [30, 31]. Gaussian Noise [32] and Gradient Clipping [33] are common techniques in Differential Privacy (DP) [34, 35] that effectively constrain the attackers from learning through the released gradients. Gradient Sparsification [36, 37] prunes the gradients through a given threshold, and Soteria [38] edits the gradients from the aspect of learned representations.
@@ II-B Neural Architecture Search (NAS)
By automatically searching the optimal model architecture, NAS algorithms have shown significant effectiveness in multiple visual tasks such as image restoration [39, 40], semantic segmentation [41, 42], and image classification [43, 44]. In image classification tasks, Zoph et al. [43] regularize the search space to a convolutional cell and construct a better architecture stacked by these cells using an RNN [45] controller. For semantic segmentation tasks, Liu et al. [42] search for the optimal network on the level of hierarchical network architecture and extend NAS to dense image prediction. In terms of image restoration tasks, Suganuma et al. [39] exploit a better Convolutional Autoencoders (CAE) with standard network components through the evolutionary search, while Chu et al. [40] discover a competitive lightweight model via both micro and macro architecture search.
Existing NAS methods adopt different strategies to exploit the search space, including evolutionary methods [46, 47, 48], Bayesian optimization [49, 50, 51], Reinforcement Learning (RL) [52, 53, 54, 43], and gradient-based search [55, 56, 57]. Different RL-based methods vary in the way to represent the agent’s policy and the optimization process. Zoph et al. [43] utilize RNN networks to sequentially encode the neural architecture and train them with the REINFORCE policy gradient algorithm [58]. Baker et al. [52] adopt Q-learning [59, 60] to train the policy network and realize competitive model design. Notably, recent works [61, 62, 63] propose training-free NAS to mitigate the issue of huge computational expenses. Instead of training from scratch, they evaluate the searched networks by some empirically designed metrics that reflect model effectiveness. In this paper, we adopt the initial gradient matching loss as the training-free search metric and provide substantial empirical evidence that such a search metric highly correlates with the real attack performance.
@@ III Problem Formulation
Figure 2: Overview of the proposed GI-NAS attack. We leverage a two-stage strategy for private batch recovery. In the first stage, we traverse the model search space and calculate the initial gradient matching loss (i.e., our training-free search metric) of each model based on the fixed input $\mathbf{z_{0}}$. We regard the model that achieves the minimal initial loss as our best model, for its performance at the start can stand out from numerous candidates. In the second stage, we adopt the architecture of the previously found best model and optimize its excessive parameters to reconstruct the private data.
GAN-based Gradient Inversion. Nevertheless, the optimization of (3) is still limited in the pixel space. Given a well pre-trained GAN, an instinctive idea is to switch the optimization from the pixel space to the GAN latent space:
@@
where $G_{\omega}$ and $\mathbf{z}\in\mathbb{R}^{B\times l}$ are respectively the generator and the latent vector of the pre-trained GAN. By reducing the optimization space from $\mathbb{R}^{B\times H\times W\times C}$ to $\mathbb{R}^{B\times l}$, (4) overcomes the uncertainty of directly optimizing the extensive pixels and exploits the abundant prior knowledge encoded in the pre-trained GAN. Based on this, recently emerged GAN-based attacks [11, 18, 12] explore various search strategies within the pre-trained GAN to utilize its expression ability.
@@
where $G_{over}$ is the over-parameterized convolutional network with excessive parameters $\mathbf{\phi}$, and $\mathbf{z_{0}}$ is the randomly generated but fixed latent code. Note that the regularization term is omitted in (5). This is because the architecture of $G_{over}$ itself can serve as implicit regularization, for convolutional networks have been discovered to possess implicit priors that prioritize clean images rather than noise as shown in [21]. Thus, the generated images that highly resemble the ground truth images can be obtained through $\mathbf{\hat{x}^{*}}=G_{over}(\mathbf{z_{0}};\mathbf{\phi^{*}})$. However, only a changeless over-parameterized network is employed for all the attack settings in [13]. As previously shown in Figure 1, although the network is over-parameterized, the attack performance exhibits significant differences when adopting different architectures. To fill this gap, we propose to further exploit such implicit architectural priors by searching the optimal over-parameterized network $G_{opt}$ for each batch. We will discuss how to realize this in Section IV.
@@ IV Method
IV-A Threat Model
Attacker’s Aim. The fundamental aim of an attacker in our method is to reconstruct the original private training data $\mathbf{x}$ from client devices with maximum fidelity by exploiting the exchanged gradients $\mathbf{g}$ in federated learning systems.
Attacker’s Knowledge. The attacker possesses three key categories of knowledge: complete access to the gradient vectors $\mathbf{g}$ transmitted during federated updates, including their structural organization and numerical values; full architectural specifications of the global model $f_{\theta}$ being trained; and the standard assumptions about input data dimensions $\mathbf{x}\in\mathbb{R}^{B\times H\times W\times C}$ without requiring distributional priors that are needed in previous GAN-based attacks [11, 12, 18]. Crucially, the knowledge requirement in our method excludes any direct access to the original training data samples or client-side information beyond the gradients, making our threat model both realistic and concerning for practical deployments.
Attacker’s Capabilities. The attacker’s capabilities encompass executing local computations to optimize reconstruction networks $G_{opt}$, deploying adaptive architectures through our NAS framework to handle varying batch characteristics, and maintaining attack efficacy even when gradients are protected by common defense mechanisms. The attacker is not allowed to modify the exchanged gradients or transmit malicious data that may compromise the training process, even if doing so can enhance the reconstruction results.
IV-B Stage 1: Training-free Optimal Model Search
Before we elaborate on the Stage 1 (i.e., Training-free Optimal Model Search), we outline its key components and objectives as follows:
  • Input for this stage: Real gradients $\mathbf{g}$, fixed latent code $\mathbf{z_{0}}$ sampled from $\mathcal{N}(0,1)$, and the model search space =$\mathcal{M}=\{G_{1},G_{2},...,G_{n}\}$.
  • Output for this stage: The searched optimal architecture $G_{opt}$ with minimal initial loss.
  • Purpose for this stage: In this stage, we traverse the model search space and calculate the initial gradient matching loss (i.e., our training-free search metric) of each model based on the fixed input $\mathbf{z_{0}}$. We regard the model that achieves the minimal initial loss as our best model $G_{opt}$.
Figure 3: The design of search space for skip connection patterns. Different skip connection patterns are determined by the skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$. $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there is not such a skip connection.
IV-B1 Model Search Space Design
@@
Search Space for Skip Connection Patterns. We assume that there are $t$ levels of encoders and decoders in total, and denote them as $e_{1},e_{2},...,e_{t}$ and $d_{1},d_{2},...,d_{t}$. As shown in Figure 3, We consider different skip connection patterns between encoders and decoders. To represent each of these patterns, we define a skip connection matrix $\mathbf{A}\in\{0,1\}^{t\times t}$ that serves as a mask to determine whether there will be new residual connections [68] between pairs of encoders and decoders. More concretely, $\mathbf{A}_{ij}=1$ indicates that there exists a skip connection from $e_{i}$ to $d_{j}$ and $\mathbf{A}_{ij}=0$ means that there is not such a skip connection. As the shapes of feature maps across different network levels can vary significantly (e.g., $64\times 64$ for the output of $e_{i}$ and $256\times 256$ for the input of $d_{j}$), we introduce connection scale factors to tackle this inconsistency and decompose all the possible scale factors into a series of 2×2$\times$2 × upsampling operations or downsampling operations with shared weights. By allocating 00 or $1$ to each of the $t^{2}$ bits in A, we broaden the search space of skip connection patterns as there are $2^{t^{2}}$ possibilities altogether and we only need to sample a portion of them.
IV-B2 Optimal Model Selection